OpenStudy (anonymous):

PLEASE HELP!!! A clerk was asked to change a $10 bill. She returned 9 more dimes than nickels and twenty-one more quarters than dimes. How many coins of each did she return? Use the V.E.S.T steps please (variable, equation, solve. try it.)

OpenStudy (anonymous):

x = nickels x + 9 = dimes (x + 9) + 21 = quarters 0.05x + 0.10(x + 9) + 0.25(x + 30) = 10.00 5x + 10(x + 9) + 25(x + 30) = 1000 5x + 10x + 90 + 25x + 750 = 1000 40x + 840 = 1000 40x = 160 x = 6 nickels x + 9 = 15 dimes x + 30 = 36 quarters

OpenStudy (anonymous):

Alright im lost howd you get the answers??? @plohrr

OpenStudy (anonymous):

if you let x = nickels x + 9 = dimes , since the clerk returned 9 more dimes than nickels , and (x+9 ) + 21 = quarters , since the clerk gave here 21 more quarters than dimes.

OpenStudy (anonymous):

yes i get that part

OpenStudy (anonymous):

ok now since a nickel is 5/100, it is 0.05. and since a dime is 10/100, it is 0.10. and quarter is 0.25. do you get it so far

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just convert the parenthesis and add them into the equation, then combine like terms until it becomes a 2 step equation.

OpenStudy (anonymous):

why did you add the 30 in tho

OpenStudy (anonymous):

21 + 9 = 30. therefore that combined is 30 into the equation lol

OpenStudy (anonymous):

oh lol okay and what I dont get is how you got the amount of coins for each one at the end.

OpenStudy (anonymous):

ok. x = 6 right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just add 9 to get 15 (# of dimes ), and add 30 to get 36 (# of quarters

OpenStudy (anonymous):

but how do you know that x=6?

OpenStudy (anonymous):

Oh nvm. i feel really stupid I get it now. lol

OpenStudy (anonymous):

haha ok